new to pre algebra

What are the correct steps to follow
7(j-5)+8=2(j+5)+5j

Answer by pwac(253)

First deal with brackets.Multipy term in bracket by number immediately preceding it.So 7(j-5) becomes 7j-35.doing this to whole equation gives:
7j-35+8=2j+10+5j we now need to collect like terms.Usual practice is to gather variables(number with letter)on left side and numbers alone on right
7j-2j-5j=10-8+35 remember change side change sign +ve becomes -ve and vice versa.
You then go on to sum both sides.However in your problem the variables
cancel each other out to leave 0 i.e
0j=37 which is an impossible equation.
Are you certain of your figures in question?
Pete